OpenStudy (anonymous):

Please help! I have a test in a few hours and I don't feel confident about some of my answers or I don't know how to answer some of the questions! If the complex number x = 3 + bi and |x|^2 = 13, which is a possible value of b?

OpenStudy (anonymous):

\[|x|^2=a^2+b^2\]

OpenStudy (anonymous):

Are there any options

OpenStudy (anonymous):

so \[|3+bi|^2=3^2+b^2=13\] solve for \(b\)

OpenStudy (anonymous):

yes, the options are \[9+b^2=13\\ b^2=4\\b=\pm2\]

OpenStudy (anonymous):

Doesnt (3+bi)^2=9+6bi-b?

OpenStudy (anonymous):

Yes these are the options: 2 4 9 10 *FYI, i might cut in and out. My school gives an IPad to each kid; instead of using them for educational purposes, kids use them to play games - taking up the WiFi in the process. So, for people like me that are trying to actually get something done: it's hard. I'll do the best i can to keep up, though.

OpenStudy (anonymous):

No, isn't it 9+6bi+bi^2. (3+bi)(3+bi) FOIL 9+3bi+3bi+bi^2 9+6bi+bi^2

OpenStudy (anonymous):

lmao schools these days .... smh @satellite73 ohhhhh is the reason why you dont do the regular process of foiling is because you only multiply the real parts with real parts and the complex values with complex values?

OpenStudy (anonymous):

no, \[|a+bi|=\sqrt{a^2+b^2}\] so \[|a+bi|^2=a^2+b^2\]

OpenStudy (anonymous):

hmmmmm confused as hell -.- Lemme google this stuff

OpenStudy (anonymous):

i though |a+bi| is the absolute value of a+bi

OpenStudy (anonymous):

Yes, I'm pretty sure that's right @Heart_Broken_Kid, unless i got something mixed up.

OpenStudy (anonymous):

hahaha sattelite is correct ... I just need to see where i went wrong